OpenStudy (anonymous):

ok solve for u -7(-4u+8)-8u = 5(u-9)-1 simplify as much as possible

OpenStudy (anonymous):

@jdoe0001

OpenStudy (caominhim):

\[-7(-4u+8)-8u = 5(u-9)-1\] first combine \[28u-56-8u = 5u-45-1\] can you take it from there?

OpenStudy (caominhim):

not combine sorry distribute

OpenStudy (anonymous):

25u-102?

OpenStudy (caominhim):

close, it looks like you forgot to switch signs

OpenStudy (caominhim):

remember that 15+3x = 5 is the same as 15-5= -3x you have to switch signs when putting the value on the other side

OpenStudy (anonymous):

so 41u - 102?

OpenStudy (caominhim):

no here\[28u - 56 - 8u = 5u-45-1\] \[28u-8u-5u=56-45-1\]

OpenStudy (anonymous):

15u = 10?

OpenStudy (caominhim):

yep, so what's u?

OpenStudy (anonymous):

15?

OpenStudy (caominhim):

no

OpenStudy (caominhim):

divide each side by 15

OpenStudy (anonymous):

0.66666666666666666667

OpenStudy (caominhim):

yea or 10/15, simplify that

OpenStudy (anonymous):

15 divided by 10 is 1.5 could that be u?

OpenStudy (caominhim):

10 over 15, not 15 over 10, you can submit an answer as .666666666666667 or you could more accurately say 2/3